# Secure Multi-Key Homomorphic Encryption with Application to Privacy-Preserving Federated Learning

### v1

@@ V-B1 Security of SMHE Construction
Simulator Construction. Let $\mathcal{I}$ be the set of $n$ honest parties, each holding a message $\mu_{i}$ and a key pair $(pk_{i},sk_{i})$ generated via MKHE.KeyGen. Let $\mathcal{C}$ be the function evaluated over the encrypted messages, and let $\mu=\mathcal{C}(\mu_{1},\dots,\mu_{n})$ be the final output.
We construct a PPT simulator $\mathsf{Sim}$ that, given only the public parameters $pp$ and the final output $\mu$, generates a view that is computationally indistinguishable from the adversary’s real-world view:
  • Encryption Simulation: SMHE employs RLWE-based encryption (e.g., CKKS or BFV). Under the RLWE assumption, the ciphertexts are semantically secure. Hence, $\mathsf{Sim}$ can simulate each ciphertext $ct_{i}$ as a uniformly random element from the ciphertext space, without knowing the corresponding plaintext $\mu_{i}$.
  • Evaluation Simulation: The evaluation phase includes ciphertext expansion, masking, and arithmetic operations. Due to the introduction of masking terms during expansion, the evaluated ciphertext $\overline{ct}$ is statistically independent of the inputs. Thus, $\mathsf{Sim}$ can simulate $\overline{ct}$ and intermediate ciphertexts using fresh samples of the appropriate algebraic structure.
@@
Security Guarantee. Let $\text{View}_{\mathcal{A}}^{\text{real}}$ denote the adversary’s view in the real execution, and let $\text{View}_{\mathsf{Sim}}^{\text{ideal}}$ be the simulated view. Under the RLWE assumption, we have:
$\text{View}_{\mathcal{A}}^{\text{real}}(\lambda)\approx_{c}\text{View}_{\mathsf{Sim}}^{\text{ideal}}(\lambda).$
That is, no efficient adversary can distinguish between the real and ideal views with non-negligible probability. This proves that SMHE achieves simulation-based security against semi-honest adversaries.
@@
V-B2 Correctness of Homomorphic Evaluation
We now present the correctness of the proposed homomorphic evaluation algorithm. We begin by analyzing the correctness in the two-party secure computation setting and subsequently extend the discussion to the general case.
Correctness of Homomorphic Addition for Two-Party Computation. The correctness of SMHE homomorphic addition is ensured by the correctness of the underlying masking scheme. Assume the two parties are $P_{1}$ and $P_{2}$, who independently generate their key materials as
$(sk_{i},pk_{i},mp_{i}=(cz_{i},\bm{\Gamma}_{i}),evk_{i})\leftarrow\texttt{SMHE.KeyGen}(pp)$
for $i=1,2$. Each party $P_{i}$ encrypts its plaintext $\mu_{i}$ into a fresh ciphertext $ct_{i}=(c_{0}^{i},c_{1}^{i})\leftarrow\texttt{SMHE.Encrypt}(\mu_{i},pk_{i}).$ The fresh ciphertexts are then expanded (associated with only the two-party encryption keys $pk_{1}$ and $pk_{2}$) as
@@ V-B2 Correctness of Homomorphic Evaluation
Decryption of $\overline{ct}_{add}$ is performed as follows: Each party $P_{i}$ computes a partial decryption
$\nu_{i}\leftarrow\texttt{SMHE.PartDec}(\bar{c}_{i},s_{i})=\bar{c}_{i}s_{i}+e_{i}\pmod{Q},i\in[1,n]$
where $e_{i}\leftarrow\chi$. The full decryption result is
[…]
=\ = $\bar{c}_{0}+\bar{c}_{1}s_{1}+\bar{c}_{2}s_{2}+e_{1}+e_{2}$
[…]
$\ c_{1}^{1}+x_{1}^{1}+z_{1}^{2},c_{1}^{2}+x_{1}^{2}+z_{1}^{1})\rangle$
=\ = $\langle sk_{1},(ct_{1}+cx_{1}+cz_{2})\rangle+\langle sk_{2},(ct_{2}+cx_{2}+cz_{1})\rangle$
[…]
Eq. (V-B2) holds due to the correctness of the masking scheme, which ensures that $\langle sk_{1},(cx_{1}+cz_{2})\rangle+\langle sk_{2},(cx_{2}+cz_{1})\rangle\ (\text{mod }Q)\approx 0$. Therefore, the correctness of the homomorphic addition for two parties in SMHE is achieved.
Correctness of Homomorphic Addition in a General Case. Let $\overline{ct}=(c_{i})_{i\in[0,n]}$ and $\overline{ct}^{\prime}=(c_{i}^{\prime})_{i\in[0,n]}$ represent two expanded ciphertexts associated with encryption keys $\{pk_{j}\}_{j\in T}$ and $\{pk_{j}\}_{j\in T^{\prime}}$, respectively. Denote the plaintexts corresponding to these ciphertexts are $\mu$ and $\mu^{\prime}$, respectively. That is, $\langle\overline{sk},\overline{ct}\rangle\approx\mu\pmod{Q}$ and $\langle\overline{sk},\overline{ct}^{\prime}\rangle\approx\mu^{\prime}\pmod{Q}$ hold, where $\overline{sk}=(1,s_{1},\cdots,s_{n})$. The masking parameters associated with the two expanded ciphertexts are given as $\left\{\left(cz_{k},\bm{\Gamma}_{k}\right)\right\}_{k\in T\cup T^{\prime}}$. The result of the homomorphic addition between the two expanded ciphertexts is $\overline{ct}_{add}=(\bar{c}_{k})_{k\in[0,n]}\leftarrow\texttt{SMHE.Add}\big{(}\overline{ct},\overline{ct}^{\prime};\{pk_{j}\}_{j\in T},\{pk_{j}\}_{j\in T^{\prime}};\{cz_{k},\bm{\Gamma}_{k}\}_{k\in T\cup T^{\prime}}\big{)}$ and its decryption proceeds as follows:
$\quad\langle\overline{sk},\overline{ct}_{add}\rangle=\bar{c}_{0}+\sum_{i\in[1,n]}s_{i}\cdot\bar{c}_{i}$ (11)
$=\big{(}(c_{0}+c_{0}^{\prime})+\sum_{i\in[1,n]}s_{i}\cdot(c_{i}+c_{i}^{\prime})\big{)}+\sum_{i\in T\cup T^{\prime}}\left\langle sk_{i},\left(cx_{i}+cz_{i}^{\prime}\right)\right\rangle$
@@
$\approx\langle\overline{sk},\overline{ct}\rangle+\langle\overline{sk},\overline{ct}^{\prime}\rangle\approx\mu+\mu^{\prime}\pmod{Q}$
@@
Eq. (11) holds due to the correctness of the masking scheme, which ensures that $\sum_{i\in T\cup T^{\prime}}\left\langle sk_{i},\left(cx_{i}+cz_{i}^{\prime}\right)\right\rangle\ (\text{mod}\ Q)\approx 0$. Therefore, the correctness of the homomorphic addition in SMHE is achieved.
@@
Correctness of Homomorphic Multiplication. The correctness of homomorphic multiplication in SMHE is ensured by the correctness of ciphertext multiplications proposed in [20] and [21]. While this applies to the multiplication of two expanded ciphertexts derived from fresh ciphertexts, we now demonstrate that the correctness also holds for multiplications between an expanded ciphertext and a masked ciphertext, as well as for multiplications between two masked ciphertexts.
@@
  • Correctness of multiplications between an expansion ciphertext and a masked ciphertext: Let $\widehat{ct}=(\hat{c}_{i})_{i\in[0,n]}$ denote a masked ciphertext, for example, $\widehat{ct}=\texttt{SMHE.Add}\big{(}\overline{ct},\overline{ct}^{\prime};\{pk_{j}\}_{j\in T},\{pk_{j}\}_{j\in T^{\prime}};\{cz_{k},\bm{\Gamma}_{k}\}_{k\in T\cup T^{\prime}}\big{)}$, with a plaintext value of $\mu+\mu^{\prime}\pmod{Q}$. Here, $\overline{ct}=(c_{i})_{i\in[0,n]}$ and $\overline{ct}^{\prime}=(c_{i}^{\prime})_{i\in[0,n]}$ are two expanded ciphertexts defined as earlier. The result of the homomorphic multiplication between the masked ciphertext $\widehat{ct}$ and the expanded ciphertext $\overline{ct}^{\prime}$ is $\overline{ct}_{\text{mult}}=$ $(\bar{c}_{i})_{i\in[0,n]}\leftarrow\texttt{SMHE.Mult}(\widehat{ct},\overline{ct}^{\prime};\{evk_{i}\}_{i\in[1,n]})$, where
@@
    $\bar{c}_{0}=$ $\hat{c}_{0}\cdot c_{0}^{\prime}+\sum_{i\in[1,n]}\big{(}\sum_{j\in[1,n]}\hat{c}_{i}\cdot c_{j}^{\prime}\boxdot\mathbf{b}_{j}\big{)}\boxdot\mathbf{v}_{i}\ (\text{mod}\ {Q}),$
    $\bar{c}_{i}=$ $\hat{c}_{i}\cdot c_{0}^{\prime}+\hat{c}_{0}\cdot c_{i}^{\prime}+\sum_{j\in[1,n]}\hat{c}_{j}\cdot c_{i}^{\prime}\boxdot\mathbf{d}_{j}$
    $+\big{(}\sum_{j\in[1,n]}\hat{c}_{i}\cdot c_{j}^{\prime}\boxdot\mathbf{b}_{j}\big{)}\boxdot\mathbf{u}_{i}\pmod{Q}$
@@
    $\quad\langle\overline{sk},\overline{ct}_{\text{mult}}\rangle=\bar{c}_{0}+\sum_{i\in[1,n]}s_{i}\cdot\bar{c}_{i}$
    $=\hat{c}_{0}c_{0}^{\prime}+\sum_{i\in[1,n]}s_{i}(\hat{c}_{i}\cdot c_{0}^{\prime}+\hat{c}_{0}\cdot c_{i}^{\prime})+\sum_{i,j=1}^{n}\big{(}(\hat{c}_{i}c_{j}^{\prime}\boxdot\mathbf{d}_{i})\cdot s_{j}$
    $\quad+\hat{c}_{i}c_{j}^{\prime}\boxdot\mathbf{b}_{j}(\mathbf{v}_{i}+s_{i}\cdot\mathbf{u}_{i})\big{)}\pmod{Q}.$
@@
    Thus, the decryption simplifies to (we define $s_{0}=1$):
    $\quad\langle\overline{sk},\overline{ct}_{\text{mult}}\rangle\approx\sum_{i\in[0,n]}\sum_{j\in[0,n]}\hat{c}_{i}c_{j}^{\prime}\cdot s_{i}s_{j}$
    $=\langle\overline{sk},\widehat{ct}\rangle\times\langle\overline{sk},\overline{ct}^{\prime}\rangle=(\mu+\mu^{\prime})\times\mu^{\prime}\pmod{Q}.$
@@
Figure 2: High-level overview of robust and privacy-preserving federated learning.
VI Application to Privacy-Preserving FL
VI-A Specific Phase of PPFL Model
We utilize the proposed SMHE to construct PPFL model. As presented in Fig. 2, PPFL contains the following steps:

### v2

@@ 5.2.1 Security of SMHE
Simulator Construction. Consider a set of $n$ parties ($n\geq 2$), each holding a message $\mu_{i}$ and a key pair $(pk_{i},sk_{i})$ generated via MKHE.KeyGen. Let $\mathcal{C}$ be the function evaluated over the encrypted messages, and let $\mu=\mathcal{C}(\mu_{1},\dots,\mu_{n})$ be the final output. We construct a PPT simulator $\mathsf{Sim}$ that, given only the public parameters $pp$ and the final output $\mu$, generates a view that is computationally indistinguishable from the adversary’s real-world view:
  • Encryption Simulation: SMHE employs RLWE-based encryption (e.g., CKKS or BFV) to generate ciphertexts (including masking parameters). Under the RLWE assumption, the ciphertexts are semantically secure. Hence, $\mathsf{Sim}$ can simulate each ciphertext $ct_{i}$ as a uniformly random element from the ciphertext space, without knowing its plaintext $\mu_{i}$.
  • Evaluation Simulation: The evaluation phase includes ciphertext expansion, masking, and arithmetic operations. Due to the introduction of masking terms, the evaluated ciphertext $\overline{ct}$ is statistically independent of the inputs. Thus, $\mathsf{Sim}$ can simulate $\overline{ct}$ and intermediate ciphertexts using fresh samples of the appropriate algebraic structure.
@@
Security Guarantee. Let $\text{View}_{\mathcal{A}}^{\text{real}}$ denote the adversary’s view in the real execution, and let $\text{View}_{\mathsf{Sim}}^{\text{ideal}}$ be the simulated view. Under the RLWE assumption, we have $\text{View}_{\mathcal{A}}^{\text{real}}(\lambda)\approx_{c}\text{View}_{\mathsf{Sim}}^{\text{ideal}}(\lambda).$ That is, no efficient adversary can distinguish between the real and ideal views with non-negligible probability. This proves that SMHE achieves simulation-based security against semi-honest adversaries.
@@
5.2.2 Correctness of Homomorphic Evaluation
We now present the correctness of the proposed homomorphic evaluation algorithm.
Correctness of Two-Party Homomorphic Addition ($\texttt{MKHE.Add}_{2}$). The correctness of SMHE homomorphic addition is ensured by the correctness of the underlying masking scheme. Assume the two parties are $P_{1}$ and $P_{2}$, who independently generate their key materials as $(sk_{i},pk_{i},evk_{i})\leftarrow\texttt{SMHE.KeyGen}(pp)$ for $i=1,2$. Each party $P_{i}$ encrypts its plaintext $\mu_{i}$ as $ct_{i}=(c_{0}^{i},c_{1}^{i})$ and generates its masking parameters $(cz_{i},\bm{\Gamma}_{i})$ by invoking $\{ct_{i},cx_{i},\bm{\Gamma}_{i}\}\leftarrow\texttt{SMHE.Encrypt}(\mu_{i},pk_{i})$. Then, $ct_{1},ct_{2}$ are expanded as
@@ 5.2.2 Correctness of Homomorphic Evaluation
Decryption of $\overline{ct}_{add}$ is performed as follows:
Each party $P_{i}$ computes a partial decryption $\nu_{i}\leftarrow\texttt{SMHE.PartDec}(\bar{c}_{i},s_{i})$ $\bar{c}_{i}s_{i}+e_{i}\pmod{Q}$, where $e_{i}\leftarrow\chi$. Then, the full decryption result is
$\mu\$$:=\texttt{SMHE.Merge}(\bar{c}_{0},\{\nu_{i}\}_{i\in[1,n]})$
$=\$${\color[rgb]{1,0,0}\lfloor(t/Q)\cdot}\big(\bar{c}_{0}+\bar{c}_{1}s_{1}+\bar{c}_{2}s_{2}+e_{1}+e_{2}\big){\color[rgb]{1,0,0}\rceil}$
$\approx\$${\color[rgb]{1,0,0}\lfloor(t/Q)\cdot}\langle(1,s_{1},s_{2}),(c_{0}^{1}+c_{0}^{2}+x_{0}^{1}+z_{0}^{2}+x_{0}^{2}+z_{0}^{1},$ (9)
$\ c_{1}^{1}+x_{1}^{1}+z_{1}^{2},c_{1}^{2}+x_{1}^{2}+z_{1}^{1})\rangle{\color[rgb]{1,0,0}\rceil}$
$=\$${\color[rgb]{1,0,0}\lfloor(t/Q)\cdot}\big(\langle sk_{1},ct_{1}+cx_{1}+cz_{2}\rangle+\langle sk_{2},ct_{2}+cx_{2}+cz_{1}\rangle\big){\color[rgb]{1,0,0}\rceil}$
@@
$\approx\$${\color[rgb]{1,0,0}\lfloor(t/Q)\cdot}\big(\langle sk_{1},ct_{1}\rangle+\langle sk_{2},ct_{2}\rangle\big){\color[rgb]{1,0,0}\rceil}\approx\mu_{1}+\mu_{2}\pmod{Q}$.
@@
Eq. (5.2.2) holds due to the correctness of the masking scheme, which ensures that $\langle sk_{1},(cx_{1}+cz_{2})\rangle+\langle sk_{2},(cx_{2}+cz_{1})\rangle\ (\text{mod }Q)\approx 0$. Therefore, the correctness of the homomorphic addition for two parties in SMHE is achieved.
Correctness of Generalized Multi-Party Homomorphic Addition (MKHE.Add). Let $\overline{ct}=(c_{i})_{i\in[0,n]}$ and $\overline{ct}^{\prime}=(c_{i}^{\prime})_{i\in[0,n]}$ represent two expanded ciphertexts associated with encryption keys $\{pk_{j}\}_{j\in T}$ and $\{pk_{j}\}_{j\in T^{\prime}}$, respectively. Denote the plaintexts corresponding to these ciphertexts are $\mu$ and $\mu^{\prime}$, respectively. That is, ${\color[rgb]{1,0,0}\lfloor(t/Q)\cdot}\langle\overline{sk},\overline{ct}\rangle{\color[rgb]{1,0,0}\rceil}\approx\mu\pmod{Q}$ and ${\color[rgb]{1,0,0}\lfloor(t/Q)\cdot}\langle\overline{sk},\overline{ct}^{\prime}\rangle{\color[rgb]{1,0,0}\rceil}\approx\mu^{\prime}\pmod{Q}$ hold, where $\overline{sk}=(1,s_{1},\cdots,s_{n})$. The masking parameters associated with the two ciphertexts are given as $\left\{\left(cz_{k},\bm{\Gamma}_{k}\right)\right\}_{k\in T\cup T^{\prime}}$. The result of the homomorphic addition between the two ciphertexts is $\overline{ct}_{add}=(\bar{c}_{k})_{k\in[0,n]}\leftarrow\texttt{SMHE.Add}\big(\overline{ct},\overline{ct}^{\prime};\{pk_{j}\}_{j\in T},\{pk_{j}\}_{j\in T^{\prime}};\{cz_{k},\bm{\Gamma}_{k}\}_{k\in T\cup T^{\prime}}\big)$ and its decryption proceeds as follows:
$\qquad{\color[rgb]{1,0,0}\lfloor(t/Q)\cdot}\langle\overline{sk},\overline{ct}_{add}\rangle{\color[rgb]{1,0,0}\rceil}=\bar{c}_{0}+\sum_{i\in[1,n]}s_{i}\cdot\bar{c}_{i}$
$={\color[rgb]{1,0,0}\lfloor(t/Q)\cdot}\big((c_{0}+c_{0}^{\prime})+\sum_{i\in[1,n]}s_{i}\cdot(c_{i}+c_{i}^{\prime})\big){\color[rgb]{1,0,0}\rceil}$ (10)
$\qquad+{\color[rgb]{1,0,0}\lfloor(t/Q)\cdot}\sum_{i\in T\cup T^{\prime}}\left\langle sk_{i},\left(cx_{i}+cz_{i}^{\prime}\right)\right\rangle{\color[rgb]{1,0,0}\rceil}$
$\approx{\color[rgb]{1,0,0}\lfloor(t/Q)\cdot}\big(\langle\overline{sk},\overline{ct}\rangle+\langle\overline{sk},\overline{ct}^{\prime}\rangle\big){\color[rgb]{1,0,0}\rceil}\approx\mu+\mu^{\prime}\pmod{Q}$.
Eq. (5.2.2) holds due to the correctness of the masking scheme, which ensures that $\sum_{i\in T\cup T^{\prime}}\left\langle sk_{i},\left(cx_{i}+cz_{i}^{\prime}\right)\right\rangle\ (\text{mod}\ Q)\approx 0$. Therefore, the correctness of the homomorphic addition in SMHE is achieved.
@@
Correctness of Homomorphic Multiplication in a General Case (MKHE.Mult). The correctness of homomorphic multiplication in SMHE is ensured by the correctness of ciphertext multiplications proposed in [20] and [21]. While this applies to the multiplication of two expanded ciphertexts derived from fresh ciphertexts, we now demonstrate that the correctness also holds for multiplications between an expanded ciphertext and a masked ciphertext, as well as for multiplications between two masked ciphertexts.
@@
  • Correctness of multiplications between an expansion ciphertext and a masked ciphertext: Let $\widehat{ct}=(\hat{c}_{i})_{i\in[0,n]}$ denote a masked ciphertext, for example, $\widehat{ct}=\texttt{SMHE.Add}\big(\overline{ct},\overline{ct}^{\prime};$ $\{pk_{j}\}_{j\in T},\{pk_{j}\}_{j\in T^{\prime}};\{cz_{k},\bm{\Gamma}_{k}\}_{k\in T\cup T^{\prime}}\big)$, with a plaintext value of $\mu+\mu^{\prime}$. Here, $\overline{ct}=(c_{i})_{i\in[0,n]}$ and $\overline{ct}^{\prime}=(c_{i}^{\prime})_{i\in[0,n]}$ are two expanded ciphertexts defined as earlier. The result of the homomorphic multiplication between the masked ciphertext $\widehat{ct}$ and the expanded ciphertext $\overline{ct}^{\prime}$ is $\overline{ct}_{\text{mult}}=$ $(\bar{c}_{i})_{i\in[0,n]}\leftarrow\texttt{SMHE.Mult}(\widehat{ct},\overline{ct}^{\prime};\{evk_{i}\}_{i\in[1,n]})$, where
@@
    $\bar{c}_{0}=\$${\color[rgb]{1,0,0}\lfloor(t/Q)\cdot}(\hat{c}_{0}\cdot c_{0}^{\prime}+\!\!\sum_{i\in[1,n]}(\!\!\sum_{j\in[1,n]}\!\!\hat{c}_{i}\cdot c_{j}^{\prime}\boxdot\mathbf{b}_{j})\boxdot\mathbf{v}_{i}){\color[rgb]{1,0,0}\rceil}(\text{mod}\ {Q}),$
    $\bar{c}_{i}=\$${\color[rgb]{1,0,0}\lfloor(t/Q)\cdot}\big(\hat{c}_{i}\cdot c_{0}^{\prime}+\hat{c}_{0}\cdot c_{i}^{\prime}+\sum_{j\in[1,n]}\hat{c}_{j}\cdot c_{i}^{\prime}\boxdot\mathbf{d}_{j}$
    $+\big(\sum_{j\in[1,n]}\hat{c}_{i}\cdot c_{j}^{\prime}\boxdot\mathbf{b}_{j}\big)\boxdot\mathbf{u}_{i}\big){\color[rgb]{1,0,0}\rceil}\pmod{Q}$
@@
    $\qquad{\color[rgb]{1,0,0}\lfloor(t/Q)\cdot}\langle\overline{sk},\overline{ct}_{\text{mult}}\rangle{\color[rgb]{1,0,0}\rceil}={\color[rgb]{1,0,0}\lfloor(t/Q)\cdot}\big(\bar{c}_{0}+\sum_{i\in[1,n]}s_{i}\cdot\bar{c}_{i}\big){\color[rgb]{1,0,0}\rceil}$
    $={\color[rgb]{1,0,0}\big\lfloor(t/Q)\cdot\lfloor(t/Q)\cdot}\big(\hat{c}_{0}c_{0}^{\prime}+\sum_{i\in[1,n]}s_{i}(\hat{c}_{i}\cdot c_{0}^{\prime}+\hat{c}_{0}\cdot c_{i}^{\prime})$
    $+\sum_{i,j=1}^{n}(\hat{c}_{i}c_{j}^{\prime}\boxdot\mathbf{d}_{i})\cdot s_{j}+\hat{c}_{i}c_{j}^{\prime}\boxdot\mathbf{b}_{j}(\mathbf{v}_{i}+s_{i}\cdot\mathbf{u}_{i}){\color[rgb]{1,0,0}\rceil\big\rceil}\!\!\!\!\pmod{Q}$.
@@
    Thus, the decryption simplifies to (Set $s_{0}=1$):
    ${\color[rgb]{1,0,0}\lfloor(t/Q)\cdot}\langle\overline{sk},\overline{ct}_{\text{mult}}\rangle{\color[rgb]{1,0,0}\rceil}\approx{\color[rgb]{1,0,0}\big\lfloor(t/Q)\cdot\lfloor(t/Q)\cdot}\!\!\!\sum_{i\in[0,n]}\sum_{j\in[0,n]}\!\!\hat{c}_{i}c_{j}^{\prime}\cdot s_{i}s_{j}{\color[rgb]{1,0,0}\rceil\big\rceil}$
    $={\color[rgb]{1,0,0}\lfloor(t/Q)\cdot}\langle\overline{sk},\widehat{ct}\rangle{\color[rgb]{1,0,0}\rceil}\!\cdot\!{\color[rgb]{1,0,0}\lfloor(t/Q)\cdot}\langle\overline{sk},\overline{ct}^{\prime}\rangle{\color[rgb]{1,0,0}\rceil}=(\mu+\mu^{\prime})\cdot\mu^{\prime}\!\!\!\!\pmod{Q}$.
@@
Figure 2: High-level overview of privacy-preserving federated learning.
6 Application to Privacy-Preserving FL
We use the proposed SMHE to construct the PPFL model, as illustrated in Fig. 2. Next, we introduce the threat model and outline the specific phases of our PPFL model.
6.1 Threat Model
We adopt a threat model similar to the one presented in the CDKS schemes [20, 21] and their privacy-preserving distributed computing applications [27, 28]. The threat model assumes the presence of multiple participants, including clients and the server, each with specific roles and assumptions regarding their behavior.
6.1.1 Participants and Assumptions
  • Clients: Each client in the federated learning system holds private data and performs local training on their dataset. We assume that clients are honest-but-curious: they faithfully follow the protocol and participate in the FL process. However, they may be curious about the data and model updates from other clients, potentially attempting to infer private information.
  • Server: The server is responsible for aggregating model updates received from clients and generating the global model. Similar to clients, the server is assumed to be honest-but-curious, faithfully executing the protocol but potentially interested in learning about the data and updates from individual clients.
6.1.2 Privacy and Security Goals
The main objective of the threat model is to ensure that sensitive information remains protected during the FL process. Specifically, the goal is to prevent the leakage of client data to other participants. The model must ensure that no participant can learn anything about others’ private data beyond what is revealed by the output of the computation. This aligns with the security objectives of MPC, where input privacy is a fundamental goal, ensuring that private data remains confidential even during collaborative computations.
6.2 Specific Phase of the PPFL Model
As depicted in Fig. 2, the PPFL model involves the following steps: